High-Low Method, Cost Formulas The controller of the South Charleston plant of Ravinia, Inc., monitored activities associated with materials handling costs. The high and low levels of resource usage occurred in September and March for three different resources associated with materials handling. The number of moves is the driver. The total costs of the three resources and the activity output, as measured by moves for the two different levels, are presented as follows: Resource Number of Moves Total Cost Forklift depreciation: Low 5,000 $1,600 High 15,000 1,600 Indirect labor: Low 5,000 $74,000 High 15,000 136,000 Fuel and oil for forklift: Low 5,000 $3,550 High 15,000 10,650 Required: If required, round your answers to two decimal places. Enter a "0" if required. 1. Determine the cost behavior formula of each resource. Use the high-low method to assess the fixed and variable components. Forklift depreciation: V $ F $ Y $ Indirect labor: V $ F $ Y $ + $ X Fuel and oil for forklift: V $ F $ Y $ X 2. Using your knowledge of cost behavior, predict the cost of each item for an activity output level of 8,000 moves. Forklift depreciation $ Indirect labor $ Fuel and oil for forklift $ 3. Construct a cost formula that can be used to predict the total cost of the three resources combined. If required, round your answers to two decimal places. Materials handling cost = $ + $ X Using this formula, predict the total materials handling cost if activity output is 8,000 moves. Y = $.
Answer #1

Working:

Number of Moves Forklift Depreciation $ Indirect
Labor $
Fuel and Oil for Forklift $
High activity level 15000 1600 136000 10650
Low activity level 5000 1600 74000 3550
Change 10000 0 62000 7100
Variable cost per unit $ 0 6.20 0.71
($62000/10000) ($7100/10000)
Fixed cost (at high level) $ 1600 43000 0
$1600 - $0 $136000 - (15000 x $6.20) $10650 - (15000 x $0.71)
Fixed cost (at low level) $ 1600 43000 0
$1600 - $0 $74000 - (5000 x $6.20) $3550 - (5000 x $0.71)

1.

Forklift Depreciation:
V $0
F $1600
Y $1600
Indirect Labor:
V $6.20
F $43000
Y $43000 + $6.20 X
Fuel and Oil for Forklift:
V $0.71
F $0
Y $0.71 X

2.

Forklift Depreciation = $1600

Indirect Labor = $43000 + ($6.20 x 8000) = $43000 + $49600 = $92600

Fuel and Oil for Forklift = $0.71 x 8000 = $5680

3.

Materials handling cost = ($1600 + $43000) + ($6.20 + $0.71) X = $44600 + $6.91 X

Y = $44600 + ($6.91 x 8000) = $44600 + $55280 = $99880